# Temporally Compressed 3D Gaussian Splatting for Dynamic Scenes

### v1

@@ 3.2.2 Gradient-Aware Mix-Precision Quantization
The influence of the Gaussian parameters on reconstruction quality is highly variable; a small adjustment in one parameter can significantly alter the rendered image, whereas similar adjustments in other parameters or in the same parameter of a different Gaussian may have minimal impact. Our approach uses gradient-based sensitivities to dynamically assign bit precision to each parameter, based on its influence on reconstruction accuracy. By leveraging this adaptive, in-optimization quantization, each parameter adjusts its quantization scale [10] in real-time, preserving detail in the reconstructed scene.
@@
We first calculate the mean sensitivity for each parameter based on the gradients [43, 32], reflecting each parameter’s impact on the image reconstruction performance. We introduce a sensitivity coefficient $S(\theta)$, which represents the responsiveness of image quality $Q$ to changes in a Gaussian parameter $\theta$. This coefficient is formulated as
$S(\theta)$ $=\frac{1}{\sum_{k=1}^{K}N_{k}}\sum_{k=1}^{K}\big{|}\frac{\partial Q_{k}}{\partial\theta}\big{|}.$ (7)
@@
where $K$ is the total number of training images used for reconstruction, $N_{k}$ is the number of pixels in the $k^{th}$ image, and $Q_{m}$ denotes the cumulative pixel intensity across the RGB channels in image $k$, serving as a proxy for image quality.
The coefficient $S(\theta)$ quantifies how sensitive $Q$ is to variations in $\theta$, measured by the gradient $\frac{\partial Q}{\partial\theta}$. A higher gradient indicates that small adjustments in $\theta$ yield large changes in image quality, suggesting a greater impact of $\theta$ on the reconstruction.
@@
By using this impact-based metric, we can effectively rank the parameters by their importance on image fidelity. We then normalize each sensitivity coefficient by scaling it based on the minimum and maximum co-efficient across all parameters. This standardization ensures that all the coefficients fall within a consistent range.
Using the normalized sensitivity $\gamma\in[0,1]$, we assign a dynamic bit precision $b$ within a range of bits $b\in[b_{\text{min}},b_{\text{max}}]$, where $b\in\mathbb{Z}$ is calculated as
$b=\left\lfloor\gamma\cdot(b_{\text{max}}-b_{\text{min}})\right\rfloor+b_{\text{min}}.$ (8)
This approach allocates higher bit precision to more sensitive parameters and lower bit precision to less sensitive ones, optimizing the balance between computational efficiency and model accuracy.
@@
After running our scene reconstruction process for a specified number of iterations, we apply mixed-precision quantization to all Gaussian parameters, excluding the position parameter, to achieve low-bit precision for the other parameters. Instead of relying on traditional vector quantization (VQ) or basic min-max quantization, we propose a parameter quantization technique with learnable scaling factors [10], integrating it directly into the optimization process rather than treating it as a post-optimization fine-tuning step. Consequently, many parameters can be effectively quantized to even 4-bit precision, reducing memory and computational load without compromising reconstruction quality.
@@
Training. We train the Gaussian parameters to model the scene one time frame at a time. Following [26], we use physically-based priors to regularize the Gaussians. The optimization objective is defined as
@@
where $\mathcal{L}_{original}$ is the same loss function as in [26], while $\mathcal{L}_{mask}$ and $\mathcal{L}_{mc}$ are defined in Eq. 4 and 5, respectively.
Algorithm 1 Keypoint Selection
@@ 3.2.3 Keypoint Interpolation
Previous implementations of dynamic 3DGS often model the time-dependence of parameters by fitting polynomials to the Gaussian parameters, which greatly limits the range of motions that can accurately be modeled. By contrast, Dynamic 3DGS [26] opts for inefficiently storing all time-dependent parameters, such as Gaussian means, rotations, and colors, for all time frames. While this greatly increases the expressivity of this method compared to other works, it comes at a significant cost in memory.
We take a different approach and observe that only a small subset of keypoints are required to accurately reconstruct complex motions. However, the placement of these keypoints across time cannot be predetermined, as it depends on the individual Gaussians. For instance, background Gaussians require a single keypoint to cover the whole sequence, whereas moving objects will need substantially more keypoints. This motivates the development of our keypoint selection strategy, which we adapt from the Ramer–Douglas–Peucker (RDP) algorithm [7]. It is applied as a post-processing step to further reduce storage requirements for the Gaussian parameters that change over time in dynamic scenes.
While the RDP algorithm selects keypoints from a sequence based on a local error tolerance, $\xi$, we propose a novel keypoint selection method, as outlined in Algorithm 1 . This method offers several advantages over the RDP algorithm. It provides greater flexibility by allowing control via both an acceptable tolerance value, $\tau$, and a maximum number of keypoints. The parameter $\tau$ defines the maximum allowable Mean Squared Error (MSE) over the sequence. Unlike RDP, which is solely controlled by $\xi$, our method enables a hard maximum bound on the number of keypoints, allowing for more precise, fine-grained control.
Following this, we flatten and transpose the time-dependent parameters from $\mathbb{R}^{T\times N\times D}$ to $\mathbb{R}^{ND\times T}$, transforming the data into $ND$ sequences of length $T$. We then compute the keypoints for all Gaussians in parallel, forming sparse matrices. The sparsity of these matrices is controlled using the parameters $max_{kp}$ and $\tau$, and and we store these sparse matrices to minimize memory usage. Upon loading the scene data, we interpolate the keypoints across all timesteps to reconstruct the dense matrices, which are then reshaped to their original parameter dimensions.
@@ 4 Experiments
We compare our method with multiple techniques extending the 3D Gaussians approach to dynamic scenes. We evaluate the methods on diverse datasets covering different real-world and synthetic scenarios. Results on additional datasets are provided in the supplementary material along with more visualizations.
4.1 Datasets
4.1.1 Panoptic Sports Dataset
We evaluate our method on the Panoptic Sports dataset, a subset from the Panoptic Studio dataset [18]. This dataset contains 6 different scenes, each having 31 camera sequences spanning 150 frames. We use 4 cameras (0, 10, 15 and 30) for testing and the rest for training, following the convention set by [26]. In addition to the images, we use the provided foreground/background segmentation to apply a segmentation loss to improve the results and prevent the background from moving. This dataset contains complex motions with objects moving quickly and over long distances.
Table 1: Quantitative results on the Panoptic dataset. The best result is shown in bold, while the second-best result is underlined. The rendering resolution is set to 640$\times$460.
Model PSNR (dB)↑ SSIM↑ LPIPS↓ FPS ↑ Storage (MB)↓
Dynamic 3DGS [26] 28.7 0.91 0.17 760 1994
STG [23] 20.5 0.78 0.44 429 19
4D Gaussian [38] 27.2 0.91 0.18 40 62
TC3DGS (Ours) 28.3 0.91 0.17 890 49
4.1.2 Neural 3D Video Dataset
The Neural 3D Video dataset [22] consists of 6 scenes with a number of cameras ranging from 18 to 21 and sequences of 300 frames. The images in this dataset are or resolution 2704$\times$2028, but we downsample them to 1352$\times$1014 for fair comparison with other methods. We hold out camera 0 for testing and do not apply any background loss, as background segmentations are not available for this dataset.
4.2 Implementation Details
For our experiments, we closely follow the hyperparameters outlined in Dynamic 3DGS [26]. Specifically, for the compression strategy, we set both the mask weight parameter $\lambda_{\text{mask}}$ and the masking consistency loss parameter $\lambda_{\text{mask-cons}}$ to $0.01$ in our primary experiments. In terms of quantization, we use Gaussian parameters with bin sizes of $b_{\text{min}}=4$ and $b_{\text{max}}=8$, while positional data is quantized to 16-bit precision to ensure spatial accuracy. For the learnable quantization step size, we use a $0.02$ learning rate. Quantization is applied after 6000 iterations in the first scene for all experiments. For keypoint interpolation, we set the tolerance to $\tau=1\times 10^{-5}$ for the Panoptic dataset and $\tau=1\times 10^{-7}$ otherwise. Moreover, we restrict the maximum number of keypoints to $\text{max}_{\text{kp}}=30$ for the Panoptic dataset and $\text{max}_{\text{kp}}=60$ otherwise, which provides an effective balance between compression and temporal trajectory accuracy. Additional ablation studies with varying values of certain hyperparameters are provided in the supplementary material.
@@ 4.3 Results
Table 2: Quantitative comparisons with various competitive baselines on the Neural 3D Video dataset. represent methods trained with a 50-frame video sequence, requiring six models to complete the evaluation.“Storage” refers to the total model size for 300 frames.
[…]
4DGS [41] 31.57 96.69 3128.00MB
4D Gaussian [38] 31.15 30.00 90MB
C-D3DGS [19] 30.46 118.00 338.00MB
Deformable 3DGS [40] 30.98 29.62 32.64MB
E-D3DGS [2] 31.20 69.70 40.20MB
STG [23] 32.04 273.47 175.35MB
Dynamic 3DGS [26] 30.97 460.00 2772.00MB
TC3DGS (Ours) 30.58 596.32 51.34MB
(a) Ground-Truth
(b) TC3DGS Render
Figure 6: Qualitative Evaluation on the Neural 3D Video dataset. This figure shows a comparison between the ground truth and our reconstruction, demonstrating that our method achieves near-identical results (PSNR: 32.7) despite extreme compression.
@@
27.9 189K 46 MB 750 32.7 59 K 53 MB 583
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Dynamic 3DGS [26] Pruned Quantized + Pruned Quantized + Pruned + KPI
Figure 7: Ablation study. Adding the different components of our method comes at virtually no cost in visual accuracy but at significantly decreased storage and rendering cost.
As shown in Table 1, our method achieves results comparable to Dynamic 3DGS [26], while utilizing, on average, 40 times less storage. Similarly, Table 2 presents the results of our method on the DyNeRF dataset, demonstrating competitive performance with a significantly smaller storage footprint. Furthermore, our approach delivers the fastest rendering speed.
On the Panoptic dataset, we failed to obtain reasonable results for the STG baseline [23], while 4D Gaussian [38] produce very distorted images, as shown in Figure 2. This dataset demonstrates the strengths of explicit methods such as Dynamic 3DGS [26] and ours. While Dynamic 3DGS [26] is able to model the complex motion and perform well on the Panoptic dataset, it is highly memory inefficient. In comparison, our model performs equally well while using much less storage.
@@ 4.4 Ablation Studies
We conduct an ablation study to evaluate the effectiveness of the individual components of our method. We report the results of this ablation in Table 3 on two scenes from the Panoptic dataset [18] and the Neural 3D Video dataset [22].
These results show the importance of each component of our method in reducing the storage size of the 3DGS [26]. In these experiments, our pruning strategy divides the number of Gaussians by 2 and by 5 folds for each scene. Then, sensitivity-aware quantization reduces the necessary storage space by 5, and keypoint interpolation by an additional 4 to 5 fold. Altogether, TC3DGS yields a compression ratio of 49 and 64, respectively. Most notably, this drastic size reduction comes at little to no cost in novel view synthesis, with an insignificant drop in PSNR of at most 0.4, which is barely perceptible.
5 Limitations
6 Conclusion
We introduced Temporally Compressed 3D Gaussian Splatting (TC3DGS), a novel framework designed to achieve memory-efficient and high-speed reconstruction of dynamic scenes. Our method achieves up to 67x compression and up to three times faster rendering while retaining high fidelity in complex motion scenarios. Through selective temporal pruning and gradient-based quantization, TC3DGS minimizes memory usage with minimal impact on visual quality. While our method is effective, limitations remain in aggressively compressing position data and handling new objects that enter the scene mid-sequence. Future work will focus on bridging the gap between spatio-temporal methods and storage-efficient dynamic scene reconstruction, aiming to extend adaptability and further reduce storage requirements for complex, evolving environments.
[…]
Supplementary Material
7 Technicolor Dataset [34]
To further demonstrate the effectiveness of our method, we evaluate our method on the Technicolor Dataset. The results in Table 4 show that our method provides competitive performance with a significantly smaller storage footprint. Figure 11 shows a comparison between the ground truth and our reconstruction, demonstrating that our method achieves near identical results.
Table 4: Per-Scene Results on the Technicolor Dataset. Our method achieves performance comparable to Dynamic3DGS [26] while significantly reducing the model size.
Method Avg. Birthday Fabien Painter Theater Train
[…]
STG [23] $33.60$ $32.09$ $35.70$ $36.44$ $30.99$ $32.58$
D. 3DGS [26] $32.12$ $30.82$ $33.62$ $34.73$ $31.12$ $30.67$
Ours $31.9$ $30.33$ $33.15$ $34.62$ $31.01$ $30.57$
[…]
STG [23] - - - - - -
D. 3DGS [26] $0.88$ $0.90$ $0.88$ $0.87$ $0.86$ $0.88$
Ours $0.85$ $0.88$ $0.86$ $0.81$ $0.84$ $0.86$
[…]
STG [23] $0.08$ $0.04$ $0.11$ $0.10$ $0.13$ $0.04$
D. 3DGS [26] $0.13$ $0.07$ $0.16$ $0.16$ $0.14$ $0.11$
Ours $0.14$ $0.09$ $0.15$ $0.17$ $0.16$ $0.13$
Table 5: Per-scene quantitative comparisons on the Panoptic Sports Dataset [18].
Method Avg. Basket. Juggle Boxes Softball Tennis Football
[…]
4DGaussians [38] $27.2$ $26.68$ $27.47$ $27.02$ $27.44$ $27.31$ $27.39$
STG [23] $20.45$ $21.60$ $19.93$ $20.65$ $19.44$ $20.82$ $20.23$
D. 3DGS [26] 28.70 28.22 29.48 29.46 28.43 28.11 28.49
Ours $27.81$ $27.92$ $29.15$ $28.28$ $27.96$ $25.97$ $28.00$
[…]
4DGaussians [38] 0.91 $0.90$ 0.92 0.91 0.92 0.92 0.92
STG [23] $0.79$ $0.78$ $0.78$ $0.79$ $0.79$ $0.78$ $0.78$
D. 3DGS [26] 0.91 0.91 0.92 0.91 $0.91$ $0.91$ $0.91$
Ours $0.89$ $0.89$ $0.90$ $0.89$ $0.89$ $0.89$ $0.89$
[…]
4DGaussians [38] $0.11$ 0.14 0.10 0.10 0.10 0.11 0.10
STG [23] 0.10 0.14 0.10 0.10 0.10 0.11 0.10
D. 3DGS [26] $0.17$ $0.18$ $0.15$ $0.17$ $0.19$ $0.18$ $0.17$
Ours $0.20$ $0.20$ $0.19$ $0.19$ $0.20$ $0.20$ $0.20$
SIZE(MB)$\downarrow$
4DGaussians [38] $63$ $66$ $59$ $63$ $57$ $72$ $65$
STG [23] 19 19 19 20 19 22 19
D. 3DGS [26] $2008$ $2161$ $1935$ $2021$ $2021$ $1915$ $2000$
Ours $49$ $46$ $32$ $44$ $41$ $48$ $34$
8 Detailed Results for Panoptic and Neural 3D Video Dataset
We present detailed results for the two datasets discussed in Sec. 4.1 of our paper, summarized in Table 5 and 8. These results include various metrics evaluated for each scene in both datasets, offering a granular view of performance. Additionally, we provide supplementary visualizations for the Neural 3D Video dataset, which further demonstrate the effectiveness of our method in capturing dynamic scenes with high fidelity. Finally, we present an ablation study over the main elements of our method in Figure 8.
9 Additional Experiments
We conducted several experiments to better understand hyperparameters of our method.
9.1 Masking and Mask Consistency
(a) SSIM $\uparrow$
Mask Consistency Loss
Mask Loss 0 0.01 0.05 0.1
0.005 0.9087 0.9105 0.9109 0.9068
0.01 0.9126 0.9140 0.9113 0.9092
0.05 0.9036 0.9006 0.8986 0.8970
0.1 0.8938 0.8956 0.8949 0.8972
(b) PSNR $\uparrow$
Mask Consistency Loss
Mask Loss 0 0.01 0.05 0.1
0.005 28.0 27.8 28.2 26.8
0.01 28.1 28.3 26.8 26.9
0.05 27.2 26.6 26.0 26.0
0.1 27.1 27.4 26.6 26.7
(c) LPIPS $\downarrow$
Mask Consistency Loss
Mask Loss 0 0.01 0.05 0.1
0.005 0.1797 0.1769 0.1781 0.1815
0.01 0.1746 0.1761 0.1748 0.1829
0.05 0.2046 0.2090 0.2139 0.2172
0.1 0.2345 0.2294 0.2344 0.2335
(d) Number of Gaussians $\downarrow$
Mask Consistency Loss
Mask Loss 0 0.01 0.05 0.1
0.005 208403 155720 143112 140897
0.01 153757 112909 98017 93568
0.05 49770 47110 29629 26154
0.1 28972 31067 19472 15656
9.2 Quantization

### v2

@@ 4.1.2 Gradient-Aware Mix-Precision Quantization
The influence of the Gaussian parameters on reconstruction quality is highly variable; a small adjustment in one parameter can significantly alter the rendered image, whereas similar adjustments in other parameters or in the same parameter of a different Gaussian may have minimal impact. Our approach uses gradient-based sensitivities to dynamically assign bit precision to each parameter, based on its influence on reconstruction accuracy. By leveraging this adaptive, in-optimization quantization, each parameter adjusts its quantization scale [Esser2020LEARNED] in real-time, preserving detail in the reconstructed scene.
@@
We first calculate the mean sensitivity for each parameter based on the gradients [Zhang_2024, Niedermayr_2024_CVPR], reflecting each parameter’s contribution on the image reconstruction performance. We introduce a sensitivity coefficient, formulated as
$\vskip-8.5359ptS(\theta)$ $=\frac{1}{\sum_{k=1}^{K}N_{k}}\sum_{k=1}^{K}\big|\frac{\partial Q_{k}}{\partial\theta}\big|.$ (7)
@@
where $K$ is the number of training images used for reconstruction, $N_{k}$ is the number of pixels in the $k^{th}$ image, and $Q_{k}$ denotes the cumulative pixel intensity across the RGB channels in image $k$.
The coefficient $S(\theta)$ quantifies the average change in pixel intensity from training views when parameter $\theta$ is changed by a small amount.
@@
By using this impact-based metric, we can effectively rank the parameters by their importance on image fidelity. We then normalize each sensitivity coefficient by scaling it based on the minimum and maximum co-efficient across all parameters. This standardization ensures that all the coefficients fall within a consistent range. Therefore, we allocate higher bit precision to more sensitive parameters and lower bit precision to less sensitive ones, optimizing the balance between computational efficiency and model accuracy.
@@
After running our scene reconstruction process for a specified number of iterations, we apply mixed-precision quantization to all Gaussian parameters, excluding the position parameter, to achieve low-bit precision for the other parameters. Instead of relying on traditional vector quantization (VQ) or basic min-max quantization, we propose a parameter quantization technique with learnable scaling factors [Esser2020LEARNED], integrating it directly into the optimization process rather than treating it as a post-optimization fine-tuning step. Consequently, many parameters can be effectively quantized to 4-bit precision, reducing memory and computational load without compromising reconstruction quality.
@@
Training. We train the Gaussian parameters to model the scene one time frame at a time. Following [luiten2023dynamic], we use physically-based priors to regularize the Gaussians. The optimization objective is defined as
@@
where $\mathcal{L}_{original}$ is the same loss function as in [luiten2023dynamic], while $\mathcal{L}_{mask}$ and $\mathcal{L}_{mc}$ are defined in Eq. 4 and 5, respectively.
Algorithm 1: Keypoint Selection
@@ 4.1.3 Keypoint Interpolation
Dynamic 3DGS [luiten2023dynamic] opts for inefficiently storing all time-dependent parameters, such as Gaussian means, rotations, and colors, for all time frames. While this greatly increases the expressivity of this method compared to other works, it comes at a significant cost in memory.
We take a different approach and observe that only a small subset of keypoints are required to accurately reconstruct complex motions. Instead of storing each gaussian’s position and rotation per frame, we identify keypoints where the trajectory of a gaussian changes significantly and interpolate linearly between them. However, the placement of these keypoints across time cannot be predetermined, as it depends on the individual Gaussians. For instance, background Gaussians require a single keypoint for the entire sequence, whereas moving objects will need substantially more keypoints. This motivates the development of our keypoint selection strategy, which we adapt from the Ramer–Douglas–Peucker (RDP) algorithm [RDP]. It is applied as a post-processing step to further reduce storage requirements for the time-varying Gaussian parameters in dynamic scenes.
While the RDP algorithm selects keypoints from a sequence based on a local error tolerance, $\xi$, we propose a novel keypoint selection method, as outlined in Algorithm 1. It provides greater flexibility by allowing control via both an acceptable tolerance value, $\tau$, and a maximum number of keypoints $max_{kp}$. The parameter $\tau$ defines the maximum allowable Mean Squared Error (MSE) over the sequence. Unlike RDP, which is solely controlled by $\xi$, our method enables a hard maximum bound on the number of keypoints, allowing for more precise, fine-grained control. In Figure 4, we illustrate with an example how our method achieves lower MSE despite storing fewer keypoints.
Following this, we flatten and transpose the time-dependent parameters from $\mathbb{R}^{T\times N\times D}$ to $\mathbb{R}^{ND\times T}$, transforming the data into $ND$ sequences of length $T$. We then compute the keypoints for all Gaussians in parallel, forming sparse matrices. The sparsity of these matrices is controlled using the parameters $max_{kp}$ and $\tau$, and and we store these sparse matrices to minimize memory usage.
@@ 5 Experiments
We evaluate the methods on diverse datasets covering different real-world scenarios. Results on additional datasets, along with dataset and implementation details, are provided in the supplementary material, which also includes visualizations. Moreover, ablation studies comparing our quantization strategy to uniform quantization are also included.
Figure 5: Our method lies on the Pareto frontier, achieving competitive performance with significantly smaller model size.
@@ 5.1 Results
As shown in Table 1 and Figure 1, our method achieves comparable results to Dynamic 3DGS [luiten2023dynamic], while using, on average, 40 times less storage. Similarly, the results on the Neural 3D Video dataset in Table 1 demonstrate competitive performance with a significantly smaller storage footprint and fastest rendering speed. These improvements are achieved without a substantial increase in training time compared to Dynamic 3DGS [luiten2023dynamic], as our method trains fewer Gaussians and just one additional parameter per Gaussian for masking.
Method Panoptic Dataset Neural 3D Video Dataset
PSNR ↑ FPS ↑ Storage ↓ PSNR ↑ FPS ↑ Storage ↓
STG 20.5 429 19MB 32.04 273 175MB
4D Gaussian 27.2 40 62MB 31.15 30 90MB
E-D3DGS 17.3 31 130MB 31.20 69 40MB
Deform. 3DGS 14.3 26 192.6MB 30.97 29 33MB
Dynamic 3DGS 28.7 760 1994MB 31.38 460 2772MB
HiCoM 26.6 456 71MB 31.17 247 270MB
TC3DGS (Ours) 28.3 890 49MB 30.96 596 51MB
Table 1: Quantitative comparisons on the Panoptic and Neural 3D Video datasets. The best result is shown in bold, and the second-best is underlined. STG on the Neural 3D Video dataset is trained on 50-frame sequences and requires six models for evaluation.
On the Panoptic dataset, we were unable to obtain reasonable results for the STG baseline [stg], while 4D Gaussian [4dgs] produce very distorted images, as shown in Figure 1. This dataset demonstrates the strengths of explicit methods such as Dynamic 3DGS [luiten2023dynamic] and ours. While Dynamic 3DGS [luiten2023dynamic] effectively models complex motion and performs well on the Panoptic dataset, it is highly memory inefficient. In contrast, our model achieves similar performance with significantly lower storage. Figure 5 shows a comparison with other methods.
@@
27.9 189K 44 MB 750 32.7 59 K 53 MB 583
@@ 5.2 Ablation Studies
We conduct an ablation study to evaluate the effectiveness of each component in our method. Table 2 presents results on two scenes from the Panoptic dataset [Joo_2019] and Neural 3D Video dataset [li2022neural3dvideosynthesis], demonstrating how each step contributes to reducing storage. Our pruning strategy reduces the number of Gaussians by 2 and by 5 folds, followed by sensitivity-aware quantization, which compresses storage by 5 times, and keypoint interpolation, adding a further  5 times reduction. Overall, TC3DGS achieves compression ratios of 49 and 64, respectively, with minimal impact on novel view synthesis, which is barely perceptible. Additional ablations are provided in the Supplementary.
5.3 Mixed-Precision Quantization